# Feasibility of Learning Weighted Automata on a Semiring

### v4

@@ 3.2. Guessable and strongly guessable functions
The definition of weakly guessable functions ensures the existence of some finite sets $Q,T\subseteq\Sigma^{*}$ and coefficients $\underline{\lambda}\in\Lambda_{Q,T}$ (writing here and from now on simply $Q$ in place of the set of rows of the Hankel matrix indexed by the words in $Q$, as discussed in Notation 2.6) that together determine a correct hypothesis automaton. In other words, functions that are not weakly guessable are those for which all hypothesis automata are incorrect, and so (in fairness to the learner!) one should immediately exclude all such functions from consideration if one seeks to learn by membership and equivalence in the style of Angluin. For weakly guessable functions, the problem then becomes: how to find a suitable $Q,T$ and $\lambda$. Clearly $Q$ must be a row-generating set for the full Hankel matrix; but is it always possible to find such a set from finitely many membership and equivalence queries? Also, even if a suitable row generating set $Q$ is found, there is no guarantee that accompanying $T$ and $\lambda$ can be found in finite time. Guessable functions are those for which there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. Strongly guessable functions are those for which for every finite row-generating set $Q$ there exists a finite subset $T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. In the following definition, the reader should recall from Remark 2.7 that the sets $\Lambda_{Q,T}$ depend upon the function $f$ (and hence also the semiring $S$).
Definition 3.11.
@@
It is clear from the definitions that every strongly guessable function is guessable, and every guessable function is weakly guessable. Moreover, observe that, by Theorem 2.12, if the pair $(Q,T)$ witnesses that $f$ is guessable, then all hypothesis automata over $Q$ and $T$ compute $f$. We now give a weaker variant of another property given in [36] to characterise guessable functions.
Definition 3.12.
@@
Remark 3.14.
As in Remarks 2.14 and 3.8, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions of co-guessable, strongly co-guessable: see Appendix A for full details. With these definitions, one can prove the following entirely analogous result.
Proposition 3.15.
@@ 3.3. Algorithmic strategies
The hierarchy we propose is (in general) orthogonal to any complexity classes - in particular there is no guarantee that a strongly guessable function can be learnt in polynomial time. See Remark 3.20 at the end of the section.
Remark 3.16.
@@
Proof.
For a weakly guessable function there exists a non-empty finite subset $Q\subseteq\Sigma^{*}$ such that $\Lambda_{Q}\neq\emptyset$. Since $\Sigma^{*}$ is countable, so is the set of non-empty finite subsets of $\Sigma^{*}$. Let $(Q_{i})_{i\in\mathbb{N}}$ be an enumeration of the finite non-empty subsets of $\Sigma^{*}$ and consider the following strategy for the learner: if $h$ is empty, query $(Q_{0},\{\varepsilon\})$; if the last query of $h$ is $(Q_{i},\{\varepsilon\})$ and does not halt the game then query $(Q_{i+1},\{\varepsilon\})$ (no matter the automata and counter-examples given by the teacher). The learner will eventually query $(Q,\{\varepsilon\})$ and since $\Lambda_{Q}\subseteq\Lambda_{Q,\{\varepsilon\}}$ the teacher (playing as an ally) will choose to declare a correct hypothesis automaton. Conversely, if the learner has a winning strategy, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and so by definition the function is weakly guessable. ∎
Proposition 3.18.
@@
Conversely, if the learner has a winning strategy in the game where the teacher plays as an adversary, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and $\Lambda_{Q,T}\setminus\Lambda_{Q}=\emptyset$ (otherwise the teacher can always choose to provide an incorrect automaton). So by definition the function is guessable. ∎
Proposition 3.19.
@@
Since $f$ is weakly guessable and for all non negative integers $i$, $Q_{i}\subseteq Q_{i+1}$ then there exists a non negative integer $I$ such that for all $i<I$, $\Lambda_{Q_{i}}$ is empty and $\Lambda_{Q_{I}}$ is non-empty. Hence by strong guessability, for all $i<I$, there exists $T$ such that $\Lambda_{Q_{i},T}$ is empty and there exists $T^{\prime}$ such that $\Lambda_{Q_{I},T^{\prime}}=\Lambda_{Q_{I}}$ is non-empty. By Remark 2.8 and Lemma 2.9, it follows that for all $i<I$, there exists $\ell$ such that for all $j\geq\ell$, $\Lambda_{Q_{i},T_{j}}$ is empty and $J$ such that for all $j\geq J$, $\Lambda_{Q_{I},T_{j}}=\Lambda_{Q_{I}}$ is non-empty (and none of the $\Lambda_{Q_{I},T}$ is empty for any $T$). Hence, the process terminates, and the startegy above is winning for the learner. ∎
Remark 3.20.
@@
  • The ascending chain condition on rows is weaker than the ascending chain condition considered in [36] (which considers all chains of submodules, including those which may not be generated by finitely many rows of the matrix) but stronger than the weak ascending chain condition from Definition 3.4 (which considers only certain chains of submodules generated by finitely many rows of the matrix). Thus a function satisfying the termination conditions above is certainly (by Theorem 3.6) weakly guessable. Weak guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo (as we have seen, it is precisely the condition required in order for a correct hypothesis automaton to exist).
@@
  • Each progress measure gives rise to a weak progress measure, and moreover, if a function has a weak progress measure then it is strongly row-bound or equivalently (by Proposition 3.13) strongly guessable. Thus the proposition above guarantees that the learning algorithm in Figure LABEL:fig:algo terminates for a class of strongly guessable functions. Strong guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo. However, it follows from the proof of Proposition 3.19 that there is an algorithm with a different (incremental) update strategy for the sets $Q$ and $T$ that will terminate with success on any strongly guessable function for which a closure strategy exists.
@@ 4.1. The general hierarchy
In other words, $\langle qa\rangle=\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes\langle p\rangle$. A similar argument shows that $\langle\varepsilon\rangle=\bigoplus_{q\in Q}\underline{\lambda}_{q}\otimes\langle q\rangle$, hence giving $\underline{\lambda}\in\Lambda_{Q}$, as required. This demonstrates that $f$ is strongly guessable. A dual argument (interchanging the rows/columns, and left/right combinations) demonstrates that $f$ is strongly co-guessable. ∎
Remark 4.2.
The proof of Lemma 4.1 actually shows that any function that is both weakly guessable and weakly co-guessable has the following stronger property. Say that a guessable function $f:\Sigma^{*}\rightarrow S$ is uniformly strongly guessable if there exists a finite set $T\subseteq\Sigma^{*}$ such that for all finite sets $Q\subseteq\Sigma^{*}$ one has $\Lambda_{Q,T}=\Lambda_{Q}$. Note the inversion of quantifiers as compared to the definition of strongly guessable; here the set $T$ has to be the same for all the sets $Q$. Dually one can define uniformly strongly co-guessable. The previous proof shows that the intersection of the class of weakly guessable and weakly co-guessable functions coincides with the intersection of the class of uniformly strongly guessable and uniformly strongly co-guessable functions. We leave open the question of whether this intersection is precisely characterised by uniform strong guessability. However, we note that it can be shown that the class of uniformly strongly guessable functions is strictly contained in the class of strongly guessable functions (as an example, the function $f_{3}$ discussed below is not uniformly strongly guessable).
@@
From this one sees that for any pair of finite subsets $Q,T\subseteq\Sigma^{*}$, the equations given in Notation 2.6 for the function $f$ (encoding certain left-linear combinations of the rows of $F$ indexed by $Q$ restricted to columns in $T$) are, due to commutativity of multiplication in $S$, equivalent to the equations encoding (right) linear combinations of the columns of $\overline{F}$ indexed by $\overline{Q}$ restricted to rows in $\overline{T}$, as obtained from A.1. Since weakly guessable, guessable and strongly guessable and their duals are described in terms of properties of these solution sets (considered for all finite sets $Q,T$), the result follows. ∎
For some semirings $S$, there are functions $f:\Sigma^{*}\rightarrow S$ that are neither weakly guessable, nor weakly co-guessable. We show that this is the case for $f_{1}$ computed by the automaton given in Figure 1(b), whether viewed over $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$, $f^{\prime}_{1}$ computed by the automaton given in Figure 1(d), whether viewed over $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$, and $f^{\prime\prime}_{1}$ computed by the automaton in Figure 1(f). There are also functions that are weakly guessable but neither guessable nor weakly co-guessable. We show that this is the case for $f_{2}$ computed by the automaton given in Figure 1(a), whether viewed as an automaton over $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$. It then follows that the mirror function $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (by Lemma 4.3). Finally, we show that the functions $f_{3}$ computed by the automaton depicted in Figure 1(a) (over $\mathbb{N}_{\max}$), $f^{\prime}_{3}$ computed by the automaton from Figure 1(c) and $f^{\prime\prime}_{3}$ computed by the automaton from Figure 1(e) are all strongly guessable but not weakly co-guessable. The functions $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable. It is in general easy to find examples at the intersection of all the classes: for commutative semirings, any function computed by an automaton that is both literal and co-literal lies in this intersection. A trivial example of such function would be the constant function $f_{4}$ mapping all words to $1_{S}$.
@@ 4.3. Collapse in some familiar semirings
In many well-studied semirings, we note that the hierarchy collapses further, due to the fact that certain ascending and descending chain conditions are satisfied. It is clear from Theorem 3.6, Theorem 3.9, Proposition 3.13, Proposition 3.15 and Lemma 4.3 that the following two properties characterise (A) when all functions computed by a finite state weighted automaton over $S$ are strongly guessable and strongly co-guessable (i.e. everything collapses, as in Figure 4(a)) and (B) when all weakly guessable functions are strongly guessable and all weakly co-guessable functions are strongly co-guessable (as in Figure 4(b)):
@@
Proof.
We give the proof for rows; the proof for columns is dual. Let $S$ be one of the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$. Let $f$ be a weakly guessable function computed over $S$, and $F$ its Hankel matrix. Let $Q\subseteq\Sigma^{*}$ be a finite set. We aim to show that there is no total $Q$-chain. First, note that, for some fixed $T$, if no row indexed by $Q$ and restricted to $T$ is fully $0_{S}$, then in each of three semirings considered, $\Lambda_{Q,T}$ is finite (possibly empty) from which it immediately follows that there is no total $Q$-chain involving $\Lambda_{Q,T}$. Let $Z_{T}=\{q\in Q\mid\langle q\rangle_{T}\mbox{ is a }0_{S}\mbox{ row}\}$, $Q^{\prime}=Q\setminus Z_{T}$ and $N_{T}\subseteq S^{Q^{\prime}\cup(Q\times\Sigma\times Q^{\prime})}$ be the set obtained from $\Lambda_{Q,T}\subseteq S^{Q\cup(Q\times\Sigma\times Q)}$ by restricting the entries $\underline{\lambda}_{p}$ and $\underline{\lambda}_{q,a,p}$ to only permit $p\in Q^{\prime}$. Clearly $Z_{T}$ is finite and its size is bounded above by $|Q|$. Notice that $N_{T}$ is also a finite set, since each $\underline{\lambda}_{p},\underline{\lambda}_{q,a,p}$ where $p\in Q^{\prime}$ can only take a finite number of different values. Moreover, it is clear from the definitions that $T\subseteq T^{\prime}$ forces $Z_{T}\supseteq Z_{T^{\prime}}$ (i.e. any row of the Hankel matrix that is zero on all columns indexed by $T^{\prime}$ is also zero on all columns indexed by $T$), $\Lambda_{Q,T}\supseteq\Lambda_{Q,T^{\prime}}$ (by Remark 2.8) and $N_{T}\supseteq N_{T^{\prime}}$ (since these sets are obtained from the previous by restricting attention to coefficients of rows indexed by $Q$ that are not identically zero on $T$ and $T^{\prime}$ respectively). Moreover, notice that if $Z_{T}=Z_{T^{\prime}}$ (the same subset of elements of $Q$ are zero when restricted to $T$ or $T^{\prime}$) and $N_{T}=N_{T^{\prime}}$ then we must have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}$. Now, suppose that $T\subsetneq T^{\prime}$ with $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$. It is clear from the above that $Z_{Q,T}\supseteq Z_{Q,T^{\prime}}$, and either this containment is strict, or if not, then $N_{Q^{\prime},T}\supsetneq N_{Q^{\prime},T^{\prime}}$. Since both sets have a finite size, there cannot be an infinite descending chain as in the definition. ∎
@@
Proof.
Suppose that $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable. Thus there is a literal automaton $\mathcal{A}$ with weights in $S^{\prime}$ that computes $f$. Viewed as a function over $S$, we also have that $f$ is weakly guessable (since $\mathcal{A}$ is a literal automaton with weights in $S$ computing $f$) and hence by assumption, strongly guessable over $S$. Suppose for contradiction that $f$ is not strongly guessable over $S^{\prime}$. Then there exists a finite set $Q$ such that for all finite sets $T$ there exists $\underline{\lambda}_{T}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$, where these solution sets are considered over $S^{\prime}$. However, noting that the coefficients $\underline{\lambda}_{T}$ lie in $S$, this directly contradicts that $f$ is strongly guessable over $S$. The final statement follows from Remark 3.10 together with the fact that any integral domain embeds into a field (its field of fractions). ∎
Remark 4.6.
@@
We now apply the results and discussion of this section to determine whether Property A or B holds in the semirings used in our examples. The table in Figure 5 demonstrates that there are interesting and important semirings where Property B holds (i.e. all weakly (co-)guessable functions are strongly (co-)guessable) whilst Property A does not hold (i.e. there exist functions that are not weakly guessable). Since over such semirings the property of being strongly guessable is characterised in terms of automata (by Theorem 3.6), it is possible for the ‘teacher’ to select a strongly guessable target function to begin with, and in light of the discussion in Remark 3.26, there is then an algorithm (similar to that of [36], but using a different update strategy) that terminates with success for all such target functions with a closure strategy.

### v5

@@ 3.2. Guessable and strongly guessable functions
The definition of weakly guessable functions ensures the existence of some finite sets $Q,T\subseteq\Sigma^{*}$ and coefficients $\underline{\lambda}\in\Lambda_{Q,T}$ (writing here and from now on simply $Q$ in place of the set of rows of the Hankel matrix indexed by the words in $Q$, as discussed in Notation 2.2) that together determine a correct hypothesis automaton. In other words, functions that are not weakly guessable are those for which all hypothesis automata are incorrect, and so (in fairness to the learner!) one should immediately exclude all such functions from consideration if one seeks to learn by membership and equivalence in the style of Angluin. For weakly guessable functions, the problem then becomes: how to find a suitable $Q,T$ and $\lambda$. Clearly $Q$ must be a row-generating set for the full Hankel matrix; but is it always possible to find such a set from finitely many membership and equivalence queries? Also, even if a suitable row generating set $Q$ is found, there is no guarantee that accompanying $T$ and $\lambda$ can be found in finite time. Guessable functions are those for which there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. Strongly guessable functions are those for which for every finite row-generating set $Q$ there exists a finite subset $T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. In the following definition, the reader should recall from Remark 2 that the sets $\Lambda_{Q,T}$ depend upon the function $f$ (and hence also the semiring $S$).
{defi}
@@
It is clear from the definitions that every strongly guessable function is guessable, and every guessable function is weakly guessable. Moreover, observe that, by Theorem 5, if the pair $(Q,T)$ witnesses that $f$ is guessable, then all hypothesis automata over $Q$ and $T$ compute $f$. We now give a weaker variant of another property given in [36] to characterise guessable functions.
{defi}
@@
Remark 18.
As in Remarks 7 and 14, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions of co-guessable, strongly co-guessable: see Appendix A for full details. With these definitions, one can prove the following entirely analogous result.
Proposition 19.
@@ 3.3. Algorithmic strategies
The hierarchy we propose is (in general) orthogonal to any complexity classes - in particular there is no guarantee that a strongly guessable function can be learnt in polynomial time. See Remark 25 at the end of the section.
Remark 20.
@@
Proof 3.3.
For a weakly guessable function there exists a non-empty finite subset $Q\subseteq\Sigma^{*}$ such that $\Lambda_{Q}\neq\emptyset$. Since $\Sigma^{*}$ is countable, so is the set of non-empty finite subsets of $\Sigma^{*}$. Let $(Q_{i})_{i\in\mathbb{N}}$ be an enumeration of the finite non-empty subsets of $\Sigma^{*}$ and consider the following strategy for the learner: if $h$ is empty, query $(Q_{0},\{\varepsilon\})$; if the last query of $h$ is $(Q_{i},\{\varepsilon\})$ and does not halt the game then query $(Q_{i+1},\{\varepsilon\})$ (no matter the automata and counter-examples given by the teacher). The learner will eventually query $(Q,\{\varepsilon\})$ and since $\Lambda_{Q}\subseteq\Lambda_{Q,\{\varepsilon\}}$ the teacher (playing as an ally) will choose to declare a correct hypothesis automaton. Conversely, if the learner has a winning strategy, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and so by definition the function is weakly guessable.
Proposition 22.
@@
Conversely, if the learner has a winning strategy in the game where the teacher plays as an adversary, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and $\Lambda_{Q,T}\setminus\Lambda_{Q}=\emptyset$ (otherwise the teacher can always choose to provide an incorrect automaton). So by definition the function is guessable.
Proposition 23.
@@
Since $f$ is weakly guessable and for all non-negative integers $i$, $Q_{i}\subseteq Q_{i+1}$ then there exists a non-negative integer $I$ such that for all $i<I$, $\Lambda_{Q_{i}}$ is empty and $\Lambda_{Q_{I}}$ is non-empty. Hence by strong guessability, for all $i<I$, there exists $T$ such that $\Lambda_{Q_{i},T}$ is empty and there exists $T^{\prime}$ such that $\Lambda_{Q_{I},T^{\prime}}=\Lambda_{Q_{I}}$ is non-empty. By Remark 3 and Lemma 4, it follows that for all $i<I$, there exists $\ell$ such that for all $j\geq\ell$, $\Lambda_{Q_{i},T_{j}}$ is empty and $J$ such that for all $j\geq J$, $\Lambda_{Q_{I},T_{j}}=\Lambda_{Q_{I}}$ is non-empty (and none of the $\Lambda_{Q_{I},T}$ is empty for any $T$). Hence, the process terminates, and the strategy above is winning for the learner.
Remark 24.
We leave open the question whether the converse of Proposition 23 holds. This is related to the question of the existence of functions that are guessable but not strongly guessable. This question remains open (see Section 4.2 and Conclusion).
Remark 25.
@@
  • The ascending chain condition on rows is weaker than the ascending chain condition considered in [36] (which considers all chains of submodules, including those which may not be generated by finitely many rows of the matrix) but stronger than the weak ascending chain condition from Definition 3.1 (which considers only certain chains of submodules generated by finitely many rows of the matrix). Thus a function satisfying the termination conditions above is certainly (by Theorem 12) weakly guessable. Weak guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo (as we have seen, it is precisely the condition required in order for a correct hypothesis automaton to exist).
@@
  • Each progress measure gives rise to a weak progress measure, and moreover, if a function has a weak progress measure then it is strongly row-bound or equivalently (by Proposition 17) strongly guessable. Thus the proposition above guarantees that the learning algorithm in Figure LABEL:fig:algo terminates for a class of strongly guessable functions. Strong guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo. However, it follows from the proof of Proposition 23 that there is an algorithm with a different (incremental) update strategy for the sets $Q$ and $T$ that will terminate with success on any strongly guessable function for which a closure strategy exists.
@@ 4.1. The general hierarchy
In other words, $\langle qa\rangle=\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes\langle p\rangle$. A similar argument shows that $\langle\varepsilon\rangle=\bigoplus_{q\in Q}\underline{\lambda}_{q}\otimes\langle q\rangle$, hence giving $\underline{\lambda}\in\Lambda_{Q}$, as required. This demonstrates that $f$ is strongly guessable. A dual argument (interchanging the rows/columns, and left/right combinations) demonstrates that $f$ is strongly co-guessable.
Remark 31.
The proof of Lemma 30 actually shows that any function that is both weakly guessable and weakly co-guessable has the following stronger property. Say that a guessable function $f:\Sigma^{*}\rightarrow S$ is uniformly strongly guessable if there exists a finite set $T\subseteq\Sigma^{*}$ such that for all finite sets $Q\subseteq\Sigma^{*}$ one has $\Lambda_{Q,T}=\Lambda_{Q}$. Note the inversion of quantifiers as compared to the definition of strongly guessable; here the set $T$ has to be the same for all the sets $Q$. Dually one can define uniformly strongly co-guessable. The previous proof shows that the intersection of the class of weakly guessable and weakly co-guessable functions coincides with the intersection of the class of uniformly strongly guessable and uniformly strongly co-guessable functions. We leave open the question of whether this intersection is precisely characterised by uniform strong guessability. However, we note that it can be shown that the class of uniformly strongly guessable functions is strictly contained in the class of strongly guessable functions (as an example, the function $f_{3}$ discussed below is not uniformly strongly guessable).
@@
From this one sees that for any pair of finite subsets $Q,T\subseteq\Sigma^{*}$, the equations given in Notation 2.2 for the function $f$ (encoding certain left-linear combinations of the rows of $F$ indexed by $Q$ restricted to columns in $T$) are, due to commutativity of multiplication in $S$, equivalent to the equations encoding (right) linear combinations of the columns of $\overline{F}$ indexed by $\overline{Q}$ restricted to rows in $\overline{T}$, as obtained from A. Since weakly guessable, guessable and strongly guessable and their duals are described in terms of properties of these solution sets (considered for all finite sets $Q,T$), the result follows.
For some semirings $S$, there are functions $f:\Sigma^{*}\rightarrow S$ that are neither weakly guessable, nor weakly co-guessable. We show that this is the case for $f_{1}$ computed by the automaton given in Figure 1(b), whether viewed over $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$, $f^{\prime}_{1}$ computed by the automaton given in Figure 1(d), whether viewed over $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$, and $f^{\prime\prime}_{1}$ computed by the automaton in Figure 1(f). There are also functions that are weakly guessable but neither guessable nor weakly co-guessable. We show that this is the case for $f_{2}$ computed by the automaton given in Figure 1(a), whether viewed as an automaton over $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$. It then follows that the mirror function $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (by Lemma 32). Finally, we show that the functions $f_{3}$ computed by the automaton depicted in Figure 1(a) (over $\mathbb{N}_{\max}$), $f^{\prime}_{3}$ computed by the automaton from Figure 1(c) and $f^{\prime\prime}_{3}$ computed by the automaton from Figure 1(e) are all strongly guessable but not weakly co-guessable. The functions $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable. It is in general easy to find examples at the intersection of all the classes: for commutative semirings, any function computed by an automaton that is both literal and co-literal lies in this intersection. A trivial example of such function would be the constant function $f_{4}$ mapping all words to $1_{S}$.
@@ 4.3. Collapse in some familiar semirings
In many well-studied semirings, we note that the hierarchy collapses further, due to the fact that certain ascending and descending chain conditions are satisfied. It is clear from Theorem 12, Theorem 15, Proposition 17, Proposition 19 and Lemma 32 that the following two properties characterise (A) when all functions computed by a finite state weighted automaton over $S$ are strongly guessable and strongly co-guessable (i.e. everything collapses, as in Figure 4(a)) and (B) when all weakly guessable functions are strongly guessable and all weakly co-guessable functions are strongly co-guessable (as in Figure 4(b)):
@@
Proof 4.3.
We give the proof for rows; the proof for columns is dual. Let $S$ be one of the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$. Let $f$ be a weakly guessable function computed over $S$, and $F$ its Hankel matrix. Let $Q\subseteq\Sigma^{*}$ be a finite set. We aim to show that there is no total $Q$-chain. First, note that, for some fixed $T$, if no row indexed by $Q$ and restricted to $T$ is fully $0_{S}$, then in each of three semirings considered, $\Lambda_{Q,T}$ is finite (possibly empty) from which it immediately follows that there is no total $Q$-chain involving $\Lambda_{Q,T}$. Let $Z_{T}=\{q\in Q\mid\langle q\rangle_{T}\mbox{ is a }0_{S}\mbox{ row}\}$, $Q^{\prime}=Q\setminus Z_{T}$ and $N_{T}\subseteq S^{Q^{\prime}\cup(Q\times\Sigma\times Q^{\prime})}$ be the set obtained from $\Lambda_{Q,T}\subseteq S^{Q\cup(Q\times\Sigma\times Q)}$ by restricting the entries $\underline{\lambda}_{p}$ and $\underline{\lambda}_{q,a,p}$ to only permit $p\in Q^{\prime}$. Clearly $Z_{T}$ is finite and its size is bounded above by $|Q|$. Notice that $N_{T}$ is also a finite set, since each $\underline{\lambda}_{p},\underline{\lambda}_{q,a,p}$ where $p\in Q^{\prime}$ can only take a finite number of different values. Moreover, it is clear from the definitions that $T\subseteq T^{\prime}$ forces $Z_{T}\supseteq Z_{T^{\prime}}$ (i.e. any row of the Hankel matrix that is zero on all columns indexed by $T^{\prime}$ is also zero on all columns indexed by $T$), $\Lambda_{Q,T}\supseteq\Lambda_{Q,T^{\prime}}$ (by Remark 3) and $N_{T}\supseteq N_{T^{\prime}}$ (since these sets are obtained from the previous by restricting attention to coefficients of rows indexed by $Q$ that are not identically zero on $T$ and $T^{\prime}$ respectively). Moreover, notice that if $Z_{T}=Z_{T^{\prime}}$ (the same subset of elements of $Q$ are zero when restricted to $T$ or $T^{\prime}$) and $N_{T}=N_{T^{\prime}}$ then we must have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}$. Now, suppose that $T\subsetneq T^{\prime}$ with $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$. It is clear from the above that $Z_{Q,T}\supseteq Z_{Q,T^{\prime}}$, and either this containment is strict, or if not, then $N_{Q^{\prime},T}\supsetneq N_{Q^{\prime},T^{\prime}}$. Since both sets have a finite size, there cannot be an infinite descending chain as in the definition.
@@
Proof 4.4.
Suppose that $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable. Thus there is a literal automaton $\mathcal{A}$ with weights in $S^{\prime}$ that computes $f$. Viewed as a function over $S$, we also have that $f$ is weakly guessable (since $\mathcal{A}$ is a literal automaton with weights in $S$ computing $f$) and hence by assumption, strongly guessable over $S$. Suppose for contradiction that $f$ is not strongly guessable over $S^{\prime}$. Then there exists a finite set $Q$ such that for all finite sets $T$ there exists $\underline{\lambda}_{T}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$, where these solution sets are considered over $S^{\prime}$. However, noting that the coefficients $\underline{\lambda}_{T}$ lie in $S$, this directly contradicts that $f$ is strongly guessable over $S$. The final statement follows from Remark 16 together with the fact that any integral domain embeds into a field (its field of fractions).
Remark 35.
@@
We now apply the results and discussion of this section to determine whether Property A or B holds in the semirings used in our examples. The table in Figure 5 demonstrates that there are interesting and important semirings where Property B holds (i.e. all weakly (co-)guessable functions are strongly (co-)guessable) whilst Property A does not hold (i.e. there exist functions that are not weakly guessable). Since over such semirings the property of being strongly guessable is characterised in terms of automata (by Theorem 12), it is possible for the ‘teacher’ to select a strongly guessable target function to begin with, and in light of the discussion in Remark 29, there is then an algorithm (similar to that of [36], but using a different update strategy) that terminates with success for all such target functions with a closure strategy.